# for the Hall C SIDIS Collaboration

### v1

@@ II.5 Pion identification
Charged pions were detected in the Super High Momentum Spectrometer S. Ali et al. , used for the first time in 2018. Considerable effort was made prior to the present experiment to understand its optical properties and acceptance, as well as to commission all of the detectors. The momentum and angle ranges used at each kinematic setting are listed in Table 1, and were chosen to provide good coverage in the region $0.3<z<0.7$, along with as much coverage in $P_{t}$ and $\phi^{*}$ as allowed by the spectrometer constraints. The polarity of the spectrometer was alternated every few days in order to separately accumulate positively and negatively charged pions. This technique provides identical acceptance for both charge states, resulting in small systematic errors in the ratios of multiplicities, compared to large acceptance devices such as CLAS CLAS Collaboration ; CLAS12 Collaboration .
As in the HMS, the SHMS detector configuration (see Fig. 6) included two pairs of segmented hodoscope planes (three scintillator planes and one quartz plane) separated by 2 m to provide fast timing signals and rough particle trajectories. The resolution in particle speed was sufficient to reject protons with momenta below 2 GeV. The average arrival time in the four paddles was compared to the arrival time of the 4 ns spaced beam micro-bunches– dubbed as the radio-frequency (RF) time. With a flight path of 22.5 m in the SHMS spectrometer, and a relative timing cut of $\pm 0.7$ ns, it was possible to remove all protons and most kaons from the event sample, as illustrated in Fig. 7. The efficiency of the timing cut was about 96% for setting III. The RF timing signal was not operational for Settings I and II.
Figure 6: Detector configuration in the SHMS
Two drift chambers, similar to those in the HMS, were used for tracking. Track pruning and detector position fiducial cuts were applied in a manner similar to the HMS. Only tracks within a fiducial volume in relative momentum ($15\%<\delta P/P<18\%$), relative in-plane angle (-30 $\theta_{y}<30 mr$), and relative out-of-plane angle (-55 $\theta_{x}<55 mr$) were kept. The tracking efficiency was found to drop from about 99.5% at low rates to about 97% at the highest rates of particles entering the detector hut. To avoid pile-up effects in the tracking, the particle rate was kept below 700 kHz by lowering the beam current to values as low as 2 $\mu A$.
@@
The heavy gas Cherenkov detector contained C4F8O at less than 1 atm pressure, giving a pion threshold of 2.61 GeV. It has four individual mirrors to focus Cherenkov light onto photomultiplier tubes. The design resulted in a small inefficient region near the center of the detector for settings I and II, and a much larger region for setting III after an unsuccessful attempt to realign the mirrors for higher efficiency, as shown in Fig. 8. Pions with momenta above 2.85 GeV were required to have tracks outside the inefficient region and a light signal greater than 1 p.e. The efficiency of this cut varied with momentum, increasing rapidly from 96% at 2.85 GeV to $99\%$ for $P_{\pi}>3.2$ GeV.
Figure 8: The $x$-position vs. $y$-position of hits on the heavy gas Cherenkov detector for setting III, showing the inefficient region that was excluded from the analysis. The color bar represents the number of photo-electrons.
@@
The noble gas Cherenkov detector was installed only for settings I and II. With a pion threshold of over 5 GeV, it was not directly used for pion identification. Its main use was to provide a clean sample of electrons for calibrating the calorimeter. Additional information on the detectors used in the experiment, plots of the trigger efficiency, and the detector efficiencies can be found in Ref. Bhatt et al. (2025); Bhatt ; Jia .
@@ II.6 Electron-pion coincidence identification
At the high luminosity of Hall C, there were many triggers for which an electron and a pion originate from different beam bunches, spaced by 4 ns for this experiment. Fig. 10 shows the electron-pion time difference distribution for a typical kinematic setting. The green lines indicate the region used to define the in-time coincidences, while the four sets of blue lines represent accidental coincidence regions—two on each side of the main peak.
Figure 10: Distribution of electron-pion time distributions for settings I and II (top panel) and setting III (bottom panel). The green vertical lines show the cuts used for true coincidences, with the four pairs of blue dashed lines showing the four regions used for accidental subtraction.
The average number of events in these four accidental peaks was subtracted from the main peak to select true electron-pion coincidences. For settings I and II (Spring 2018), the peak width was about 0.4 ns, consistent with the expected timing resolution of the trigger scintillators in both spectrometers, and a cut of $\pm 1$ ns was used to identify in-time coincidences. For setting III (Fall 2018), a mis-cabling problem caused the peak to be much wider (2.2 ns), reducing the ability to reject kaons and protons using coincidence timing. A wider cut of $\pm 2$ ns was therefore applied for setting III. Fortunately, the RF timing was operational for setting III, which more than compensated for this deficiency. The accidental-to-real ratio varied throughout the experiment from 10% to 50%.
@@ II.7 Readout trigger and data acquisition
The trigger consisted of in-time signals on any three out of the four hodoscope planes in each spectrometer. This ensured essentially 100% trigger efficiency. The timing resolution of each plane was about 0.5 ns, resulting in an accuracy of typically 0.3 ns for the electron-pion time difference, when using multiple planes. The trigger signaled the Data Acquisition (DAQ) system CEBAF Data Acquisition System () (DAQ) to read out pulse time and height information for all the detectors in both spectrometers, and record them at rates of up to 3000 Hz. In Spring 2018, a buffering system was not in place, which meant that up to 40% of events were not recorded. Buffering mode was implemented for Fall 2018, resulting in 100% of events being recorded. Data runs typically lasted one hour.
During the Spring 2018 runs (settings I and II), several problems caused a rate-dependent loss of information for the desired electron-pion coincidences. These included a non-optimal configuration of the Flash ADC modules used to read out pulse heights, non-optimal timing windows for the multi-hit TDCs, and the use of more than one trigger type, such as pre-scaled triggers from the HMS or SHMS only. The multiplicative correction factor $C_{DT}$ for these effects, determined by running with different beam currents under otherwise identical conditions, was parameterized as: $C_{DT}=1.03+0.19(R_{HMS}+R_{SHMS})$ where $R_{HMS}$ and $R_{SHMS}$ are the trigger rates in the two spectrometers, in MHz. The factor of 1.03 at zero luminosity was obtained by comparing a few measurements made at identical kinematic settings during Setting I and III, under the assumption that systematic normalization errors were under much better control after Spring 2018. After many improvements to the hardware setup during the summer of 2018, the correction was found to be much smaller: $C_{DT}=1+0.04(R_{HMS}+R_{SHMS})$.
@@ III Monte Carlo simulation
A Monte Carlo (MC) simulation Hall C MC , named SIMC, was performed for each target and pair of spectrometer settings primarily to model the spectrometer acceptance and evaluate radiative corrections. Another important use was to model pion and kaon decays (which led to muons and pions in the SHMS detector hut, respectively). For each setting, the simulation was used to generate a large number of events for three distinct physics processes: charged pion SIDIS itself (see Sec. III.1); and the two background processes, exclusive pion production and the semi-exclusive $\pi\Delta(1232)$ final-state arising from bremsstrahlung emitted by either the incoming or the outgoing electron. These radiative contributions were treated in the angle peaking approximation using the formalism of Mo and Tsai Mo and Tsai (1969). The simulation also includes a detailed model of the targets, and geometrical acceptance and magnetic field maps of the spectrometer magnets. The MC accounted for energy loss and multiple scattering in the target, vacuum windows, and detectors. Meson decays were allowed at any point along the particle trajectory, with the charged decay products tracked through the remainder of the spectrometer. The MC has been demonstrated to accurately reproduce the performance of the Hall C spectrometers Bhatt . The multiplicity and cross section models used in the simulation are described in the next three sections. The SIDIS model was improved by scaling the ratio of measured yields to the MC yield and iterating this process.
@@ III.1 SIDIS model
The inclusive DIS cross section $\sigma_{DIS}(x,Q^{2},\epsilon)$ is from a global fit to all world data available by the year 2020 for electrons scattering from both proton and deuteron targets Bosted and Christy (2008). It is the most comprehensive model of the DIS cross section measured with the electron spectrometer used in this experiment, and the inclusive data collected at the kinematic settings used in this paper were found to be consistent with this model to within a few percent.
@@
where, $M_{p/n\pi^{\pm}}(z,x,Q^{2})$ are the charged pion multiplicities from the proton (p) and neutron (n), $q_{i}$ are the quark charges, the quark distribution functions $u,d,s,\bar{u},\bar{d},\bar{s}$ were taken from CTEQ5 Kuhlmann (1999), and the favored and unfavored fragmentation functions $D_{f}$ and $D_{u}$ were parameterized as:
@@
where, $s_{v}=\ln(Q^{2}/2)$ and the target mass corrections were applied using Schienbein et al. (2008)
@@
i.e., a Gaussian distribution with the parameter $b=(0.12z^{2}+0.2)^{-1}$ GeV-2, common to all processes with amplitude $M_{0}(z,x,Q^{2})$. Note that we do not have any azimuthal dependence in this fit, consistent with the results of the present experiment. Also note that we do not have a factorized expression: the multiplicity function depends on the electron variables $(x,Q^{2},W)$, which we found necessary to describe the data of this experiment.
@@ III.2 Exclusive pion production model
The parameters $p_{i}$ for exclusive pion production from the proton ($ep\rightarrow e\pi^{+}n$) and the neutron ($en\rightarrow e\pi^{-}p$) are obtained from fits to world data on LT separated pion electroproduction cross sections Horn and are shown in Tables 4 and  5 .
@@ III.6 Event selection cuts and efficiency corrections
The quality of the models and corrections used in the simulation is demonstrated in Fig. 13, showing the excellent agreement between the experimental yields and simulated yields for setting I with the SHMS spectrometer centered on $z=0.9$ to capture the contributions for exclusive pion production (centered on electron-pion missing mass $M_{x}=0.94$ GeV), $\pi\Delta$ electroproduction (centered on $M_{x}=1.232$ GeV), and high-$z$ SIDIS from both the target liquid and endcap (“dummy target”).
@@ IV.1 Data analysis
The results discussed in this paper included the additional cut $M_{x}>1.6$ GeV, to remove the region where contributions from nucleon resonances, semi-exclusive processes, and higher-twist effects appear to be large, as shown in Fig. 2. This cut was removed in the version of the analysis used to iterate the SIDIS model used in the MC simulation.
Numerical results for the multiplicities are tabulated in a full three-dimensional grid in $(z,P_{t},\phi^{*})$ for each target, pion polarity, and HMS setting in $(x,Q^{2})$ on the Hall C experimental results web page Hall C pion-SIDIS results . In this table, each HMS spectrometer setting was divided in two, with relative scattering angle either positive or negative. A total of 20,000 bins are listed, based on the criterion that the Monte Carlo simulation prediction was for more than 4 counts, to ensure approximately Gaussian statistical errors on the experimental data. The table also includes results from thirteen additional HMS settings taken in Fall 2018 and Spring 2019 to study charge-symmetry violation in pion fragmentation functions, as reported in Ref. Bhatt et al. (2025). These settings covered a small range $\left<P_{t}\right$$\sim$ 0.1 GeV, and therefore are not included in the results of the present publication. The tables also include multiplicity results with no radiative corrections applied, which may prove useful in future global fits with consistent radiative correction models and formalism.
@@ IV.2 Pion multiplicities as a function of $(z,P_{t},\phi^{*})$
Figure 14: Pion multiplicity as a function of $\phi^{*}$ for $x=0.31$, $Q^{2}=3$ GeV2 (kinematic setting I) for eight bins in $P_{t}$ (left to right) and four target/final-state configurations (top to bottom), for four values of $z$ as indicated on the right edge of the rightmost panels. The solid curves are fits to each dataset at fixed $z$, $P_{t}$, target, and pion charge with the functional form $M_{0}[1+A\cos(\phi^{*})+B\cos(2\phi^{*})]$.
Figure 15: Same as Fig. 14 except for $x=0.30$, $Q^{2}=4.1$ GeV2 (kinematic setting II).
Figure 16: Same as Fig. 14 except for $x=0.45$, $Q^{2}=4.5$ GeV2 (kinematic setting III).
@@
In terms of the standard structure functions Bacchetta et al. (2007),
@@ IV.3 Pion multiplicities averaged over $\phi^{*}$
Figure 17: The $\phi^{*}$ averaged fit parameter $M_{0}$ weighted by $z^{2}$, as a function of $P_{t}$, for the three kinematic settings of this experiment (top to bottom) in four bins in $z$ (left to right). Within each panel, the results from top to bottom are for $\pi^{+}$ from a proton target (black), $\pi^{+}$ from a deuteron target (red), $\pi^{-}$ from a deuteron target (green) and $\pi^{-}$ from a proton target (blue). The curves are the predictions of the MAPS collaboration based on a fit to previous world data Bacchetta et al. (2022, 2024); Cerutti , normalized with a parameter $k$ to give the best overall agreement with these results. The dashed curves illustrate two-parameter Gaussian fits to each dataset of the form $M=ae^{(-bP_{t}^{2})}$.
The results for the $\phi^{*}$ averaged parameter, $M_{0}$, from the fits described above are displayed in Fig. 17 as a function of $P_{t}$ for the three kinematic settings, the target and pion charge combinations, in four bins in $z$. The measured multiplicities are compared to the calculation of MAPS Bacchetta et al. (2022, 2024); Cerutti scaled by $P_{t}$-independent normalization factors $k$ that give the best agreement with these data. The MAPS calculations, which use a combination of Gaussian and weighted Gaussian distributions in $P_{t}$, based on a fit to data from HERMES and COMPASS, are generally in good agreement with the measured $P_{t}$-dependence.
Figure 18: Normalization factors $k$ that best match our results and the MAPS calculations Bacchetta et al. (2022, 2024); Cerutti for the three kinematic settings of this experiment (top to bottom), as a function of $z$, and the four flavor cases (color scheme same as in Fig. 17).
@@
The $P_{t}$-independent normalization factors $k$ are plotted in Fig. 18 as a function of $z$, for the target and pion charge combinations and the three kinematic settings. They are, on average, closest to unity for setting II ($W=3.3$ GeV), and tend to be larger than unity for setting I ($W=2.8$ GeV), and even larger for setting III ($W=2.6$ GeV). There is also a clear trend for $k$ to increase with increasing $z$, especially for positive pions from the proton target, and to a lesser extent for the positive pions from the deuteron target. These trends are likely related to the fact that the present data are at lower $W$ and higher $x$ than the HERMES and COMPASS data that went into the MAPS global fit.

### v2

@@ II.5 Pion identification
Charged pions were detected in the Super High Momentum Spectrometer shms_nim, used for the first time in 2018. The momentum and angle ranges used at each kinematic setting are listed in Table 1, and were chosen to provide good coverage in the region $0.3<z<0.7$, along with as much coverage in $P_{t}$ and $\phi^{*}$ as allowed by the spectrometer constraints. The polarity of the spectrometer was alternated every few days in order to separately accumulate positively and negatively charged pions. This technique provides identical acceptance for both charge states, resulting in small systematic errors in the ratios of multiplicities, compared to large acceptance devices such as CLAS CLAS6col; clas12col.
As in the HMS, the SHMS detector configuration included two pairs of segmented hodoscope planes (three scintillator planes and one quartz plane) separated by 2 m to provide fast timing signals and rough particle trajectories. The resolution in particle speed was sufficient to reject protons with momenta below 2 GeV. The average arrival time in the four paddles was compared to the arrival time of the 4 ns spaced beam micro-bunches– dubbed as the radio-frequency (RF) time. With a flight path of 22.5 m in the SHMS spectrometer, and a relative timing cut of $\pm 0.7$ ns, it was possible to remove all protons and most kaons from the event sample, as illustrated in Fig. 4. The efficiency of the timing cut was about 96% for setting III. The RF timing signal was not operational for Settings I and II.
Two drift chambers, similar to those in the HMS, were used for tracking. (The tracking efficiency was found to drop from about 99.5% at low rates to about 97% at the highest rates of particles entering the detector hut. To avoid pile-up effects in the tracking, the particle rate was kept below 700 kHz by lowering the beam current to values as low as 2 $\mu A$.
@@
The heavy gas Cherenkov detector contained C4F8O at less than 1 atm pressure, giving a pion threshold of 2.61 GeV. It has a small inefficient region near the center of the detector for settings I and II, and a much larger region for setting III. Pions with momenta above 2.85 GeV were required to have tracks outside the inefficient region and a light signal greater than 1 p.e. The efficiency of this cut varied with momentum, increasing rapidly from 96% at 2.85 GeV to $99\%$ for $P_{\pi}>3.2$ GeV.
@@
The noble gas Cherenkov detector was installed only for settings I and II. With a pion threshold of over 5 GeV, it was not directly used for pion identification. Its main use was to provide a clean sample of electrons for calibrating the calorimeter. Additional information on the detectors used in the experiment, plots of the trigger efficiency, and the detector efficiencies can be found in Ref. Bhatt:2024prq; hem_thesis; shuo_thesis.
@@ II.6 Electron-pion coincidence identification
At the high luminosity of Hall C, there were many triggers for which an electron and a pion originate from different beam bunches, spaced by 4 ns for this experiment. The average number of events in four accidental peaks (2 on either side of the true coincidence peak) was subtracted from the central peak to select true electron-pion coincidences. For settings I and II (Spring 2018), the peak width was about 0.4 ns, consistent with the expected timing resolution of the trigger scintillators in both spectrometers, and a cut of $\pm 1$ ns was used to identify in-time coincidences. For setting III (Fall 2018), a mis-cabling problem caused the peak to be much wider (2.2 ns), reducing the ability to reject kaons and protons using coincidence timing. A wider cut of $\pm 2$ ns was therefore applied for setting III. Fortunately, the RF timing was operational for setting III, which more than compensated for this deficiency. The accidental-to-real ratio varied throughout the experiment from 10% to 50%.
@@ II.7 Readout trigger and data acquisition
The trigger consisted of in-time signals on any three out of the four hodoscope planes in each spectrometer. This ensured essentially 100% trigger efficiency. The timing resolution of each plane was about 0.5 ns, resulting in an accuracy of typically 0.3 ns for the electron-pion time difference. The trigger signaled the Data Acquisition (DAQ) system coda to read out pulse time and height information for all the detectors in both spectrometers, and record them at rates of up to 3000 Hz.
During the Spring 2018 runs (settings I and II), several problems caused a rate-dependent loss of information for the desired electron-pion coincidences. The multiplicative correction factor $C_{DT}$ for these effects, determined by running with different beam currents under otherwise identical conditions, was parameterized as: $C_{DT}=1.03+0.19(R_{HMS}+R_{SHMS})$ where $R_{HMS}$ and $R_{SHMS}$ are the trigger rates in the two spectrometers, in MHz. The factor of 1.03 at zero luminosity was obtained by comparing to Spring 2018 runs (setting III), for which most of the problems of settings I and II were fixed, and the luminosity correction is much smaller: $C_{DT}=1+0.04(R_{HMS}+R_{SHMS})$.
@@ III Monte Carlo simulation
A Monte Carlo (MC) simulation simc, named SIMC, was performed for each target and pair of spectrometer settings primarily to model the spectrometer acceptance and evaluate radiative corrections. Another important use was to model pion and kaon decays (which led to muons and pions in the SHMS detector hut, respectively). For each setting, the simulation was used to generate a large number of events for three distinct physics processes: charged pion SIDIS itself (see Sec. III.1); and the two background processes, exclusive pion production and the semi-exclusive $\pi\Delta(1232)$ final-state arising from bremsstrahlung emitted by either the incoming or the outgoing electron. These radiative contributions were treated in the angle peaking approximation using the formalism of Mo and Tsai tsai. The simulation also includes a detailed model of the targets, and geometrical acceptance and magnetic field map of the spectrometer magnets. The MC accounted for energy loss and multiple scattering in the target, vacuum windows, and detectors. Meson decays were allowed at any point along the particle trajectory, with the charged decay products tracked through the remainder of the spectrometer. The MC has been demonstrated to accurately reproduce the performance of the Hall C spectrometers hem_thesis. The multiplicity and cross section models used in the simulation are described in the next three sections. The SIDIS model was improved by scaling the ratio of measured yields to the MC yield and iterating this process.
@@ III.1 SIDIS model
The inclusive DIS cross section $\sigma_{DIS}(x,Q^{2},\epsilon)$ is from a global fit to all world data available by the year 2020 for electrons scattering from both proton and deuteron targets bosted_fit. It is the most comprehensive model of the DIS cross section measured with the electron spectrometer used in this experiment, and the inclusive data collected at the kinematic settings used in this paper were found to be consistent with this model to within a few percent.
@@
where, $M_{p/n\pi^{\pm}}(z,x,Q^{2})$ are the charged pion multiplicities from the proton (p) and neutron (n), $q_{i}$ are the quark charges, the quark distribution functions $u,d,s,\bar{u},\bar{d},\bar{s}$ were taken from CTEQ5 cteq5, and the favored and unfavored fragmentation functions $D_{f}$ and $D_{u}$ were parameterized as:
@@
where, $s_{v}=\ln(Q^{2}/2)$ and the target mass corrections were applied using Schienbein_2008
@@
i.e., a Gaussian distribution with the parameter $b=(0.12z^{2}+0.2)^{-1}$ GeV-2, common to all processes with amplitude $M_{0}(z,x,Q^{2})$. Note that we do not have any azimuthal dependence in this fit, consistent with the $\pi^{+}$ results of the present experiment. Also note that we do not have a factorized expression: the multiplicity function depends on the electron variables $(x,Q^{2},W)$, which we found necessary to describe the data of this experiment.
@@ III.2 Exclusive pion production model
The parameters $p_{i}$ for exclusive pion production from the proton ($ep\rightarrow e\pi^{+}n$) and the neutron ($en\rightarrow e\pi^{-}p$) are obtained from fits to world data on LT separated pion electroproduction cross sections th_thesis and are shown in Tables 4 and  5 .
@@ III.4 Kinematic dependence of radiative corrections
The exclusive pion and $\pi\Delta$ corrections were assumed to be subtractive, and reduced the measured cross section by as much as 7% in the worst case. The multiplicative correction factor ranged from 0.8 at high $z$ to 1.2 at low $z$.
@@ III.5 Acceptance corrections
The SHMS and HMS momentum range used for this analysis was -16% <dp/p<<dp/p< +18% and -9% <dp/p<<dp/p< +11%, respectively, for central angles. For extreme values of in-plane and out-of-plane angles, for which acceptance corrections were more than $\approx$ 10% away from unity, more restrictive $dp/p$ ranges were used. The SHMS and HMS in-plane relative angle range for this analysis was -30 mr $y^{{}^{\prime}}_{tar}<$ 30 mr, while the out of plane relative angle range was -55 mr $x^{{}^{\prime}}_{tar}<$ 55 mr and -$x^{{}^{\prime}}_{tar}< 65 mr$ respectively.
@@ III.6 Event selection cuts and efficiency corrections
The quality of the models and corrections used in the simulation is demonstrated in Fig. 8, showing the good agreement between the experimental yields and simulated yields for setting I with the SHMS spectrometer centered on $z=0.9$ to capture the contributions for exclusive pion production (centered on electron-pion missing mass $M_{x}=0.94$ GeV), $\pi\Delta$ electroproduction (centered on $M_{x}=1.232$ GeV), and high-$z$ SIDIS from both the target liquid and endcap (“dummy target”).
@@ IV.1 Data analysis
The results discussed in this paper included the additional cut $M_{x}>1.6$ GeV, to remove the region where contributions from nucleon resonances, semi-exclusive processes, and higher-twist effects appear to be large, as was found in Ref. Tigran07. This cut was removed in the version of the analysis used to iterate the SIDIS model used in the Monte Carlo simulation.
Numerical results for the multiplicities are tabulated in a full three-dimensional grid in $(z,P_{t},\phi^{*})$ for each target, pion polarity, and HMS setting in $(x,Q^{2})$ on the Hall C experimental results web page webpage. In this table, each HMS spectrometer setting was divided in two, with relative scattering angle either positive or negative. A total of 20,000 bins are listed, based on the criterion that the Monte Carlo simulation prediction was for more than 4 counts, to ensure approximately Gaussian statistical errors on the experimental data. The table also includes results from thirteen additional HMS settings taken in Fall 2018 and Spring 2019 to study charge-symmetry violation in pion fragmentation functions, as reported in Ref. Bhatt:2024prq. These settings covered a small range $\left<P_{t}\right$$\sim$ 0.1 GeV, and therefore are not included in the results of the present publication. The tables also include multiplicity results with no radiative corrections applied, which may prove useful in future global fits with consistent radiative correction models and formalism.
@@ IV.2 Pion multiplicities as a function of $(z,P_{t},\phi^{*})$
Figure 9: Pion multiplicity as a function of $\phi^{*}$ for $x=0.31$, $Q^{2}=3$ GeV2 (kinematic setting I) for eight bins in $P_{t}$ (left to right) and four target/final-state configurations (top to bottom), for four values of $z$ as indicated on the right edge of the rightmost panels. Relative systematic errors (not shown) are estimated to be 2.2% point-to-point and 1.8% for the scale uncertainty. The solid curves are fits to each dataset at fixed $z$, $P_{t}$, target, and pion charge with the functional form $M_{0}[1+A\cos(\phi^{*})+B\cos(2\phi^{*})]$. No fully differential results from previously published experiments are available to compare to.
Figure 10: Pion multiplicity as a function of $\phi^{*}$ for $x=0.30$, $Q^{2}=4.1$ GeV2 (kinematic setting II) for eight bins in $P_{t}$ (left to right) and four target/final-state configurations (top to bottom), for four values of $z$ as indicated on the right edge of the rightmost panels. Relative systematic errors (not shown) are estimated to 2.2% point-to-point and 1.8% for the scale uncertainty. The solid curves are fits to each dataset at fixed $z$, $P_{t}$, target, and pion charge with the functional form $M_{0}[1+A\cos(\phi^{*})+B\cos(2\phi^{*})]$. No fully differential results from previously published experiments are available to compare to.
Figure 11: Pion multiplicity as a function of $\phi^{*}$ for $x=0.45$, $Q^{2}=4.5$ GeV2 (kinematic setting III) for eight bins in $P_{t}$ (left to right) and four target/final-state configurations (top to bottom), for four values of $z$ as indicated on the right edge of the rightmost panels. Relative systematic errors (not shown) are estimated to be 1.3% point-to-point and 1.8% for the scale uncertainty. The solid curves are fits to each dataset at fixed $z$, $P_{t}$, target, and pion charge with the functional form $M_{0}[1+A\cos(\phi^{*})+B\cos(2\phi^{*})]$. No fully differential results from previously published experiments are available to compare to.
@@
In terms of the standard structure functions Bacchetta07,
@@ IV.3 Pion multiplicities averaged over $\phi^{*}$
Figure 12: The $\phi^{*}$ averaged fit parameter $M_{0}$ weighted by $z^{2}$, as a function of $P_{t}$, for the three kinematic settings of this experiment (top to bottom) in four bins in $z$ (left to right). Within each panel, the results from top to bottom are for $\pi^{+}$ from a proton target (black), $\pi^{+}$ from a deuteron target (red), $\pi^{-}$ from a deuteron target (green) and $\pi^{-}$ from a proton target (blue). Relative systematic errors (not shown) are estimated to be 1.3% (2.2%) point-to-point for setting III (settings I and $I$) and 1.8% for the scale uncertainty. The curves are the predictions of the MAP collaboration based on a fit to previous world data MAPS; MAP24; MAP22_4, normalized with a parameter $k$ to give the best overall agreement with these results. The dashed curves illustrate two-parameter Gaussian fits to each dataset of the form $M=ae^{(-bP_{t}^{2})}$. The arrows point to the values of $P_{t}$ for which $y_{h}=-1$, where $y_{h}$ is the definition of rapidity used in Ref. Boglione:2022gpv.
The results for the $\phi^{*}$ averaged parameter, $M_{0}$, from the fits described above are displayed in Fig. 12 as a function of $P_{t}$ for the three kinematic settings, the target and pion charge combinations, in four bins in $z$. The measured multiplicities are compared to the calculation of MAP MAPS; MAP24; MAP22_4 scaled by $P_{t}$-independent normalization factors $k$ that give the best agreement with these data. The MAP calculations, which use a combination of Gaussian and weighted Gaussian distributions in $P_{t}$, based on a fit to data from HERMES and COMPASS, are generally in good agreement with the measured $P_{t}$-dependence.
Figure 13: Normalization factors $k$ that best match our results and the MAP calculations MAPS; MAP24; MAP22_4 for the three kinematic settings of this experiment (top to bottom), as a function of $z$, and the four flavor cases (color scheme same as in Fig. 12).
@@
The $P_{t}$-independent normalization factors $k$ are plotted in Fig. 13 as a function of $z$, for the target and pion charge combinations and the three kinematic settings. They are, on average, closest to unity for setting II ($W=3.3$ GeV), and tend to be larger than unity for setting I ($W=2.8$ GeV), and even larger for setting III ($W=2.6$ GeV). There is also a clear trend for $k$ to increase with increasing $z$, especially for positive pions from the proton target, and to a lesser extent for the positive pions from the deuteron target. These trends are likely related to the fact that the present data are at lower $W$ and higher $x$ than the HERMES and COMPASS data that went into the MAP global fit.